Find the remainder when x303+x302+x301+x300............+x+1 is divided by x3-x
 

Grade:9

1 Answers

Harsh Patodia IIT Roorkee
6 years ago
X3-X=0
X(X2-1)=0
X=0,1,-1
NOW IF X=0 THE given expression is 1
If X=1 the given expression is=304
If x=-1 then it will be=0